OpenStudy (anonymous):

Solve for θ if 0º ≤ θ < 360º 4 cos θ - 3 sec θ = 0

OpenStudy (kc_kennylau):

Remember that secθ is 1/cosθ :)

OpenStudy (campbell_st):

well try rewriting the equation as \[4 \cos(x) = 3 \sec(x) ... or ..... 4\cos(x) = \frac{3}{\cos(x)}\] then perhaps multiplying both sides of the equation by cos(x) to start

OpenStudy (anonymous):

i think this can be solved like this \[4\cos \theta - 3(1divcos \theta)=0\] or,\[4 \cos ^{2}\theta =3\] \[or,\cos \theta=\sqrt{3\div4}\] \[or,\cos \theta=\sqrt{3}\div2\] and now we know that cos\[\sqrt{3}\div2 is cos 30,etc\] in this way we can find the values of\[ \theta upto 360\]

OpenStudy (campbell_st):

well I would have thought \[\cos(\theta) = \pm \frac{\sqrt{3}}{2}\]